# Orthogonalized Policy Optimization Decoupling Sampling Geometry from Optimization Geometry in RLHF

### v2

@@ 7 Discussion
OPO can be viewed as a generalization of existing alignment methods:
  • Setting $\alpha=1$ and removing the quadratic term recovers SFT-like behavior.
  • Setting $\alpha\to 0$ and using KL geometry yields objectives with similar peak-seeking and exponential-type weighting effects as in KL-regularized preference optimization (e.g., DPO).
  • OPO occupies the previously unexplored region of the design space: peak-seeking sampling with quadratic regularization in ratio coordinates.
@@
The OPO objective (Equation˜7) involves expectations under $\pi_{\mathrm{ref}}$. In practice:
@@ 8 Conclusion
We have presented Orthogonalized Policy Optimization (OPO), a framework that decomposes alignment objectives into two independent design axes: sampling geometry and optimization geometry. By working in ratio coordinates where Pearson $\chi^{2}$ divergence naturally induces quadratic penalties, OPO avoids the entangled dynamics of KL-based methods and provides a well-conditioned objective with linear gradient dynamics. Our analysis shows that OPO maintains stable optimization even in high-confidence regimes, addressing a key limitation of existing methods in reasoning tasks. We believe OPO provides a principled foundation for future work on robust and scalable LLM alignment.
@@ A.1 Log-Ratio Approximation Error
Let $\Delta_{\theta}(y)=\log\pi_{\theta}(y)-\log\pi_{\mathrm{ref}}(y)$ be the log-ratio, and $v_{\theta}(y)=\exp(\Delta_{\theta}(y))-1$ be the exact ratio deviation. Let $\delta=\|\Delta_{\theta}\|_{\infty}$. For $\delta<1$ (trust region regime):
@@ A.2 Equivalence to $\chi^{2}$-Constrained Maximization
A.3 Distributional Stability Guarantees

### v3

@@ 7 Discussion
OPO can be viewed as occupying a previously unexplored point in the design space:
  • $\alpha\to 1$ (conservative escort): $\omega_{\alpha}\to 1$, recovering supervised fine-tuning–like uniform weighting.
  • $\alpha\to 0$ (aggressive escort) with KL geometry: the exponential-type weighting resembles KL-regularized preference optimization (e.g., DPO), but inherits gradient saturation.
  • OPO: moderate $\alpha$ (peak-seeking escort) combined with Euclidean ($\chi^{2}$) geometry—non-saturating gradients with advantage-aware sampling.
@@
The OPO objective (Equation˜9) involves expectations under $\pi_{\mathrm{ref}}$. In practice:
@@ 8 Conclusion
We have presented Orthogonalized Policy Optimization (OPO), a framework grounded in two independent geometries: an $\alpha$-divergence escort projection that shapes sampling, and a Euclidean mirror map (Pearson $\chi^{2}$) that shapes optimization. The same loss emerges from two complementary derivations—mirror descent and ratio-space regression—confirming it is uniquely determined by the geometry choice. The resulting objective has a constant Hessian, linear non-saturating gradients, and global contraction, addressing a key limitation of KL-based methods in high-confidence reasoning tasks. We believe OPO provides a principled foundation for future work on robust and scalable LLM alignment.
@@ A.1 Log-Ratio Approximation Error
Let $\Delta_{\theta}(y)=\log\pi_{\theta}(y)-\log\pi_{\mathrm{ref}}(y)$ be the log-ratio, and $v_{\theta}(y)=\exp(\Delta_{\theta}(y))-1$ be the exact ratio deviation. Let $\delta=\|\Delta_{\theta}\|_{\infty}$. For sufficiently small $\delta$ (in particular $\delta<1$, which is enforced by the $\chi^{2}$ trust-region penalty $\frac{\mu}{2}\mathbb{E}[v^{2}]$ bounding $\|\Delta\|_{\infty}$):
@@ A.2 Equivalence to $\chi^{2}$-Constrained Maximization
A.3 Proof of Proposition 4.2: $\alpha$-Geometric Interpolation
Proof.
Define the oracle target $P^{*}(y)\propto\exp(A(y))$ and the parameterized family $\rho_{\alpha}(y)\propto\tilde{p}(y)^{\alpha}\,P^{*}(y)^{1-\alpha}$ for $\alpha\in[0,1]$.
Step 1 (Geodesic structure). In the $e$-affine coordinate system of the statistical manifold [11], the log-density of $\rho_{\alpha}$ is:
$\log\rho_{\alpha}(y)=\alpha\log\tilde{p}(y)+(1-\alpha)\log P^{*}(y)-\log Z_{\alpha}$ (24)
where $Z_{\alpha}$ is the normalizing constant. Since $\log\rho_{\alpha}$ is an affine function of $\alpha$ in the natural parameter, $\{\rho_{\alpha}\}_{\alpha\in[0,1]}$ forms an $e$-geodesic (exponential geodesic) connecting $P^{*}$ (at $\alpha=0$) to $\tilde{p}$ (at $\alpha=1$) in the exponential family.
Step 2 (Importance weight). The sampling weight is:
$\omega_{\alpha}(y)=\frac{\rho_{\alpha}(y)}{\tilde{p}(y)}\propto\left(\frac{P^{*}(y)}{\tilde{p}(y)}\right)^{1-\alpha}$ (25)
In the on-policy setting where $P^{*}(y)=\exp(A(y))$ and $\tilde{p}=\pi_{\mathrm{old}}$, this yields $\omega_{\alpha}\propto(\exp(A)/\tilde{p})^{1-\alpha}$, which simplifies to $\exp((1-\alpha)A)$ when the $\tilde{p}$ correction is absorbed by the sampling measure.
Step 3 (Geometric consistency). This interpolation corresponds to the exponential geodesic in the statistical manifold and is consistent with Rényi $\alpha$-divergence interpolation geometry [11]. Under the dually flat structure of exponential families, it coincides with the $e$-projection onto the geodesic submanifold connecting $\tilde{p}$ and $P^{*}$. ∎
A.4 Distributional Stability Guarantees